OpenStudy (anonymous):

how do you solve a equation like this 7/8 + X=4

OpenStudy (anonymous):

To isolate a variable (f.e. X): subtract, add, multiply and/or divide to get numbers on one side and variables on the other side

OpenStudy (kamibug):

You want to solve for x, right?

OpenStudy (anonymous):

yah

OpenStudy (vivek3461):

x = 4 - (7/8)

OpenStudy (kamibug):

Okay, so like @Tajar said, you need the x alone on one side of the equation. This one is pretty simple. All you gotta do is move 7/8 to the other side. Since you're moving it to the opposite side of the equal sign, you need to make 7/8 negative. So you'd have x = 4 - 7/8. Get it? Think you can take it from there? =)

OpenStudy (anonymous):

thanks got it

OpenStudy (kamibug):

No problem! Glad to help! =) I've always loved algebraic equations! xD